OpenStudy (anonymous):

what is equal to 6rs ^{9}div42r ^{5}s ^{2}

OpenStudy (anonymous):

s^7/7r^4

OpenStudy (anonymous):

Did that help?

OpenStudy (anonymous):

yes it did

OpenStudy (anonymous):

Okay awesome if u have any other questions then go ahead and ask

OpenStudy (anonymous):

\[8-\left(\begin{matrix}7 \\ c\end{matrix}\right) \div 4+ \left(\begin{matrix}2 \\ 3c\end{matrix}\right)\]

OpenStudy (anonymous):

@sydneymmelin

OpenStudy (anonymous):

Hey im sorry i dont jnow tne answer

OpenStudy (anonymous):

Sorry know the

OpenStudy (anonymous):

ok np thanks :)